# Enhancing adversarial robustness in Natural Language Inference using explanations

### v1

@@ Abstract
The surge of state-of-the-art Transformer-based models has undoubtedly pushed the limits of NLP model performance, excelling in a variety of tasks. We cast the spotlight on the underexplored task of Natural Language Inference (NLI), since models trained on popular well-suited datasets are susceptible to adversarial attacks, allowing subtle input interventions to mislead the model. In this work, we validate the usage of natural language explanation as a model-agnostic defence strategy through extensive experimentation: only by fine-tuning a classifier on the explanation rather than premise-hypothesis inputs, robustness under various adversarial attacks is achieved in comparison to explanation-free baselines. Moreover, since there is no standard strategy of testing the semantic validity of the generated explanations, we research the correlation of widely used language generation metrics with human perception, in order for them to serve as a proxy towards robust NLI models. Our approach is resource-efficient and reproducible without significant computational limitations [^1] The code will be available upon publication..
@@ 1 Introduction
  • In order to facilitate (approximate) explanation evaluation, we provide an association between linguistic quality of explanations and model robustness, verified by humans.
@@ 4.2 Results
Attack success rate ($\downarrow$) 62.9% 59.41% 63.81% 58.88% 61.2%
@@
Table 5: Attack results synopsis for attacking P or H using TextFooler. Bold values denote best results.
@@
Table 6: Attack results synopsis for attacking P or H using BERT-Attack. Bold values denote best results.
@@
The robustness guarantees are strongly associated with the quality of the explanations themselves: the linguistic quality of explanations, as well as the human perception of correctness (Table 1) are consistently correlated with ExplainThenPredict model robustness, with RoBERTa arising as the most potent encoder, as all other components of ExplainThenPredict remain invariant. Thus, since the choice of Seq2Seq module matters, we safely conclude that optimizing explanation quality results in advanced ExplainThenPredict robustness. As a byproduct of this observation, we can state that leveraging ExplainThenPredict to advance NLI robustness is not sufficient on its own, and the weight needs to be shifted towards producing more faithful and linguistically correct explanations.

### v2

@@ Abstract
The surge of state-of-the-art transformer-based models has undoubtedly pushed the limits of NLP model performance, excelling in a variety of tasks. We cast the spotlight on the underexplored task of Natural Language Inference (NLI), since models trained on popular well-suited datasets are susceptible to adversarial attacks, allowing subtle input interventions to mislead the model. In this work, we validate the usage of natural language explanation as a model-agnostic defence strategy through extensive experimentation: only by fine-tuning a classifier on the explanation rather than premise-hypothesis inputs, robustness under various adversarial attacks is achieved in comparison to explanation-free baselines. Moreover, since there is no standard strategy for testing the semantic validity of the generated explanations, we research the correlation of widely used language generation metrics with human perception, in order for them to serve as a proxy towards robust NLI models. Our approach is resource-efficient and reproducible without significant computational limitations. [^1] The source code is publicly available in: https://github.com/alexkoulakos/explain-then-predict.
@@ 1 Introduction
  • In order to facilitate (approximate) explanation evaluation, we provide an association between metrics for linguistic quality of explanations and model robustness, as verified by humans.
@@ 4.2 Results
Attack success rate ($\downarrow$) 62.9% 59.41% 61.5% 58.88% 61.2%
@@
Table 5: Attack results synopsis for attacking P or H using TextFooler. Bold values denote best results (row-wise).
@@
Table 6: Attack results synopsis for attacking P or H using BERT-Attack. Bold values denote best results (row-wise).
@@
The robustness guarantees are strongly associated with the quality of the explanations themselves: the linguistic quality of explanations, as well as the human perception of correctness (Table 1) are consistently correlated with ExplainThenPredict model robustness, with RoBERTa arising as the most potent encoder [^8] An interesting future work would be to experiment with baseline classifier architectures other than BERT (e.g. ALBERT, DistilBERT, RoBERTa) and examine if we get similar results., as all other components of ExplainThenPredict remain invariant. Thus, since the choice of Seq2Seq module matters, we safely conclude that optimizing explanation quality results in advanced ExplainThenPredict robustness. As a byproduct of this observation, we can state that leveraging ExplainThenPredict to advance NLI robustness is not sufficient on its own, and the weight needs to be shifted towards producing more faithful and linguistically correct explanations.
@@ 5 Conclusion
Acknowledgments
The research work was supported by the Hellenic Foundation for Research and Innovation (HFRI) under the 3rd Call for HFRI PhD Fellowships (Fellowship Number 5537).